# Revisiting Graph Neural Networks on Graph-level Tasks: Comprehensive Experiments, Analysis, and Improvements

### v1

@@ 4.4.2. Efficiency Metric
We evaluate the efficiency of models on both graph classification and regression tasks based on the training time (s), inference time (s), memory usage (MB) in training and inference phrases.
Table 3. Evaluation on graph classification. The best and the second performance are highlighted in bold and in underline.
Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
SN IMDB-B Acc 0.6840 0.7100 0.6860 0.6860 0.7440 0.6980 0.7150 0.7290 0.7070 0.7250 0.7030 0.7310 0.6230 0.6970 0.7550
F1 0.6963 0.7117 0.6991 0.7105 0.7625 0.7146 0.7158 0.7391 0.7014 0.7185 0.7078 0.7422 0.6321 0.6940 0.7693
IMDB-M Acc 0.4620 0.4820 0.4650 0.4670 0.5080 0.4750 0.4727 0.4967 OOM 0.4720 0.4650 0.5067 0.4130 0.4867 0.5233
F1 0.4393 0.4706 0.4512 0.4435 0.4850 0.4563 0.4466 0.4839 OOM 0.4443 0.4376 0.4874 0.3867 0.4788 0.5087
REDDIT-B Acc 0.9305 0.8965 0.9094 0.9280 0.9195 0.9260 OOM OOM OOM 0.8276 OOM 0.8625 OOM OOM 0.9325
F1 0.9322 0.9026 0.9132 0.9284 0.9263 0.9303 OOM OOM OOM 0.8298 OOM 0.8722 OOM OOM 0.9330
COLLAB Acc 0.7644 0.7328 0.7406 0.7556 0.8164 0.7696 OOM OOM OOM 0.7528 0.6988 0.8278 OOM 0.7688 0.7862
F1 0.7454 0.7099 0.7169 0.7395 0.7973 0.7520 OOM OOM OOM 0.7114 0.6702 0.8046 OOM 0.7320 0.7432
BIO PROTEINS Acc 0.7286 0.7250 0.7367 0.7277 0.7440 0.7286 0.7061 0.7475 0.7071 0.7322 0.7304 0.7232 0.7015 0.7367 0.7457
F1 0.6596 0.6432 0.6617 0.6712 0.7058 0.6598 0.5838 0.6978 0.6026 0.6641 0.6653 0.6442 0.6386 0.6993 0.6998
DD Acc 0.7310 0.7233 0.7542 0.7156 0.7802 0.7343 OOM 0.7776 0.7335 0.7632 0.7598 0.7632 OOM OOM 0.7632
F1 0.6676 0.6555 0.6707 0.6384 0.7252 0.6687 OOM 0.7267 0.6402 0.6845 0.6841 0.6844 OOM OOM 0.6886
ENZYMES Acc 0.4650 0.4833 0.5183 0.4800 0.4950 0.4783 0.4617 0.5267 0.4650 0.4417 0.4667 0.5250 0.4917 0.5333 0.5533
F1 0.4775 0.4781 0.5098 0.4794 0.4237 0.4737 0.4592 0.5279 0.4628 0.4337 0.4695 0.5309 0.4902 0.5314 0.5671
MolHIV Acc 0.9696 0.9701 0.9695 0.9687 0.9691 0.9696 0.9689 0.9728 0.9703 0.9686 0.9689 0.9701 0.9686 0.9705 0.9742
F1 0.3192 0.3466 0.2991 0.2221 0.2861 0.3213 0.2537 0.3682 0.3413 0.2203 0.2518 0.2502 0.2225 0.3628 0.3693
MolTOX21 Acc 0.5548 0.5553 0.5529 0.5523 0.5412 0.5514 0.5463 0.5561 0.5580 0.5283 0.5350 0.5384 0.5112 0.5565 0.5727
Mi-F1 0.9120 0.9114 0.9096 0.9083 0.9109 0.9101 0.9121 0.9138 0.9137 0.8993 0.9004 0.9048 0.8986 0.9126 0.9174
Ma-F1 0.3628 0.3601 0.3418 0.2240 0.3452 0.3765 0.3858 0.3850 0.3815 0.2020 0.2125 0.2180 0.1981 0.3847 0.3882
CHE MUTAG Acc 0.8041 0.8570 0.8199 0.8094 0.8254 0.8091 0.8091 0.8409 0.8062 0.7643 0.7706 0.7790 0.7067 0.8607 0.9102
F1 0.8583 0.8907 0.8644 0.8590 0.8642 0.8506 0.8558 0.8770 0.8512 0.8255 0.8296 0.8447 0.8012 0.8912 0.9310
NCI1 Acc 0.8158 0.8154 0.8146 0.8169 0.7662 0.8160 0.7866 0.8187 0.7603 0.7816 0.7825 0.7852 0.6995 0.7560 0.8253
F1 0.8160 0.8153 0.8149 0.8171 0.7708 0.8168 0.7914 0.8190 0.7673 0.7827 0.7831 0.7855 0.6997 0.7557 0.8249
MoIBACE Acc 0.6711 0.6645 0.6477 0.6756 0.6513 0.6749 0.7125 0.6803 0.6078 0.6019 0.6689 0.6294 0.6089 0.7099 0.7895
F1 0.7154 0.7351 0.7099 0.7264 0.7366 0.7436 0.7400 0.7382 0.7055 0.7001 0.6545 0.7077 0.7047 0.7366 0.8242
MolPCBA Acc 0.5456 0.5466 0.5451 0.5461 0.5447 ¿3 days OOM 0.5523 0.5414 OOM OOM OOM OOM OOM 0.5531
Mi-F1 0.9850 0.9853 0.6945 0.9849 0.9836 ¿3 days OOM 0.9860 0.9850 OOM OOM OOM OOM OOM 0.9858
Ma-F1 0.1200 0.1500 0.1317 0.1333 0.0167 ¿3 days OOM 0.2140 0.2064 OOM OOM OOM OOM OOM 0.2152
Table 4. Evaluation on graph regression. The best and the second performance are highlighted in bold and in underline.
Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
GE 3-Cycle MAE 0.4396 0.3962 0.5119 0.4290 0.4251 0.4238 0.0192 0.0022 0.0009 0.8776 0.4369 0.5410 0.4744 0.4231 0.0006
R2 0.6970 0.7546 0.8081 0.7044 0.7107 0.5679 0.9996 0.9998 0.9999 0.0096 0.6850 0.5003 0.6062 0.7174 0.9999
4-Cycle MAE 0.2806 0.2536 0.5407 0.2772 0.2720 0.2696 0.0148 0.0223 0.0062 0.7851 0.2745 0.5442 0.5398 0.2734 0.0039
R2 0.8229 0.8864 0.4011 0.8331 0.8477 0.8406 0.9997 0.9990 0.9998 0.0157 0.8350 0.4032 0.4436 0.8606 0.9999
5-Cycle MAE 0.2782 0.1858 0.4608 0.2399 0.2338 0.2659 0.0716 0.0340 0.0115 0.9021 0.2763 0.4532 0.4570 0.2180 0.0078
R2 0.8334 0.9364 0.4815 0.8614 0.8936 0.8553 0.9862 0.9972 0.9996 0.0019 0.8190 0.6081 0.4939 0.9060 0.9998
6-Cycle MAE 0.3014 0.1895 0.4690 0.2302 0.1788 0.2926 0.0862 0.0584 0.0368 0.8876 0.3035 0.4557 0.5027 0.1773 0.0275
R2 0.7926 0.9162 0.6189 0.8803 0.9279 0.8073 0.9569 0.9956 0.9969 0.0021 0.7891 0.6219 0.6077 0.9301 0.9982
7-Cycle MAE 0.4011 0.2111 0.5898 0.2852 0.1569 0.3942 0.1560 0.0563 0.0488 0.8273 0.4220 0.5713 0.5836 0.1500 0.0414
R2 0.5887 0.8635 0.4594 0.7921 0.9354 0.6091 0.9460 0.9937 0.9946 0.0064 0.5487 0.4883 0.4692 0.9404 0.9957
8-Cycle MAE 0.4757 0.2632 0.5285 0.2910 0.1377 0.4701 0.1153 0.0493 0.0399 0.7430 0.4790 0.5340 0.4808 0.1286 0.0337
R2 0.3570 0.7220 0.2212 0.7553 0.9428 0.3840 0.9703 0.9949 0.9963 0.0330 0.3520 0.2051 0.3319 0.9466 0.9974
4-Path MAE 0.7154 0.4273 0.7340 0.5273 0.1606 0.6361 0.0243 0.0150 0.0083 0.7784 0.5019 0.7316 0.7270 0.1512 0.0066
R2 0.1574 0.6349 0.0898 0.5270 0.9459 0.3188 0.9921 0.9996 0.9998 0.0279 0.5390 0.0909 0.1310 0.9534 0.9999
5-Path MAE 0.6848 0.3949 0.6669 0.5189 0.1559 0.6133 0.0134 0.0153 0.0092 0.7510 0.4846 0.6933 0.7053 0.1407 0.0081
R2 0.1050 0.6358 0.1675 0.5015 0.9473 0.3254 0.9995 0.9995 0.9998 0.0283 0.5204 0.0823 0.0650 0.9562 0.9999
6-Path MAE 0.6164 0.3913 0.6594 0.5629 0.1394 0.6930 0.0138 0.0133 0.0089 0.7571 0.4552 0.6627 0.6589 0.1299 0.0079
R2 0.2470 0.5958 0.1286 0.4133 0.9539 0.1408 0.9995 0.9996 0.9998 0.0250 0.5459 0.1282 0.1280 0.9607 0.9998
4-Clique MAE 0.3432 0.3449 0.3498 0.3430 0.3427 0.3578 0.0091 0.0092 0.0005 0.3866 0.2355 0.3422 0.3424 0.1797 0.0004
R2 0.1612 0.1348 0.1086 0.1213 0.1022 0.1057 0.9662 0.9964 0.9999 0.0890 0.8896 0.1604 0.1610 0.9002 0.9999
Tailed Triangle MAE 0.3506 0.2886 0.3812 0.3465 0.3665 0.3405 0.0189 0.0146 0.0034 0.8934 0.3791 0.3891 0.4095 0.3277 0.0018
R2 0.7884 0.8606 0.7303 0.7813 0.7673 0.8053 0.9989 0.9995 0.9999 0.0176 0.7336 0.7008 0.6245 0.8177 0.9999
Chordal Cycle MAE 0.4377 0.3683 0.4390 0.4216 0.3690 0.4246 0.0295 0.0336 0.0044 0.8720 0.4441 0.4413 0.4728 0.3671 0.0028
R2 0.6346 0.7109 0.6002 0.6414 0.6848 0.6421 0.9970 0.9951 0.9999 0.0333 0.5746 0.5841 0.5684 0.7293 0.9999
Triangle Rectangle MAE 0.4657 0.4034 0.4794 0.4736 0.4068 0.4890 0.3619 0.3464 0.3271 0.8263 0.4286 0.4755 0.5409 0.4301 0.3123
R2 0.5211 0.6291 0.5335 0.4921 0.6235 0.4702 0.6895 0.7063 0.7241 0.0338 0.5828 0.5037 0.3852 0.5691 0.7501
@@ 4.5. Effectiveness Evaluation
In this subsection, we evaluate the effectiveness of GNNs on graph tasks from two aspects. Firstly, in Section 4.5.1, we assess whether existing GNNs can capture the core semantic patterns of graphs, where the core semantic patterns mainly determine graph labels. Secondly, in Section 4.5.2, we evaluate the expressivity ability of GNNs based, i.e., whether GNNs can distinguish graph geometric patterns, such as the number of cycles and paths.
@@
Table 5. Efficiency evaluation. We report the total training time (Train (s)), and total inference time (Infer (s)), and GPU memory during training (T-Mem) and inference (I-Mem).
Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
SN REDDIT-B Train(s) 192.44 182.98 179.55 218.03 758.22 8365.60 OOM OOM OOM OOM OOM 371.86 OOM OOM 411.93
Infer(s) 28.42 26.21 26.04 32.95 30.45 1953.91 OOM OOM OOM OOM OOM 79.35 OOM OOM 34.50
T-Mem(MB) 287 288 310 287 2120 360 OOM OOM OOM OOM OOM 604 OOM OOM 4287
I-Mem(MB) 190 122 135 190 244 194 OOM OOM OOM OOM OOM 289 OOM OOM 190
BIO ENZYMES Train(s) 26.05 26.34 25.23 31.95 38.85 375.25 312.46 350.25 470.05 120.48 67.5 75.29 495.23 57.53 55.45
Infer(s) 5.19 5.33 5.04 6.84 7.17 76.30 68.74 68.93 94.92 36.12 19.20 28.27 85.11 7.76 5.80
T-Mem(MB) 83 78 77 83 228 103 582 2369 1656 323 186 506 282 324 246
I-Mem(MB) 42 36 36 42 46 45 270 879 790 152 95 194 54 58 42
CHE MUTAG Train(s) 10.78 11.03 10.05 13.50 15.96 60.13 52.05 88.70 67.33 44.37 35.32 39.99 80.49 25.15 18.72
Infer(s) 3.86 3.89 3.51 5.27 4.01 17.74 18.91 21.18 25.10 15.80 11.63 15.18 29.78 7.64 4.22
T-Mem(MB) 42 44 44 42 82 46 422 775 1306 56 48 431 134 140 102
I-Mem(MB) 19 20 21 19 24 19 245 210 722 26 27 74 24 25 19
GE 4-Cycle Train(s) 166.50 160.95 140.88 180.57 1224.39 482.95 1180.51 645.97 1487.66 395.71 284.34 394.60 1770.94 300.21 1503.16
Infer(s) 256.17 241.02 235.40 279.43 2058.62 1033.68 1776.91 920.13 2219.63 549.60 441.95 583.55 2675.22 461.87 2255.94
T-Mem(MB) 88 85 85 88 380 104 1214 1734 3080 262 269 470 238 237 4327
I-Mem(MB) 64 62 62 64 260 68 342 608 726 104 152 270 103 96 737

### v2

@@ 4.4.2. Efficiency Metric
We evaluate the efficiency of models on both graph classification and regression tasks based on the training time (s), inference time (s), memory usage (MB) in training and inference phases.
4.5. Hyperparameter and Hardware Setting
For classification datasets, we set the batch size as 32 for four larger datasets (REDDIT, COLLAB, DD, and MolPCBA) and 128 for the other datasets. For regression datasets, we set the batch size as 256. To efficiently tune the hyperparameters for each model, we employed the Optuna framework (Akiba et al., 2019). For each model, we conducted $200$ trials using the Tree-structured Parzen Estimator (TPE) sampler and a MedianPruner to terminate unpromising trials early. The hyperparameter search spaces were defined as follows: the hidden dimension is selected from $\{64,128,256,512\}$, the learning rate from $\{1e-2,1e-3,1e-4,1e-5\}$, GNN layers from $\{1,2,3,4\}$, and the dropout rate from $\{0,0.1,0.2,0.3,0.4,0.5\}$. All models are trained for a maximum of 2000 epochs, with early stopping applied if no improvement is observed on the validation set within 50 epochs.
All experiments are executed on a CentOS 7 machine equipped with dual 10-core Intel® Xeon® Silver 4210 CPUs @ 2.20GHz, 8 NVIDIA GeForce RTX 2080 Ti GPUs (11GB each) and 256GB RAM.
@@ 4.6. Effectiveness Evaluation
Table 4. Evaluation on graph classification. The best and second performance are highlighted in bold and underlined, respectively. “Met.” means Metrics, “OOM” means out-of-memory, and “>3>3days” means training cannot be completed within 3 days.
Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
SN I-B Acc 0.6840 0.7100 0.6860 0.7280 0.6860 0.7440 0.6980 0.7150 0.7290 0.7070 0.7250 0.7030 0.7310 0.6230 0.6970 0.7120
F1 0.6963 0.7117 0.6991 0.7391 0.7105 0.7625 0.7146 0.7158 0.7391 0.7014 0.7185 0.7078 0.7422 0.6321 0.6940 0.7147
I-M Acc 0.4620 0.4820 0.4650 0.4980 0.4670 0.5080 0.4750 0.4727 0.4967 OOM 0.4720 0.4650 0.5067 0.4130 0.4867 0.4833
F1 0.4393 0.4706 0.4512 0.4842 0.4435 0.4850 0.4563 0.4466 0.4839 OOM 0.4443 0.4376 0.4874 0.3867 0.4788 0.4726
RED Acc 0.9029 0.8965 0.9094 OOM 0.9280 0.9195 0.9260 OOM OOM OOM 0.8276 OOM 0.8625 OOM OOM OOM
F1 0.9123 0.9026 0.9132 OOM 0.9322 0.9263 0.9303 OOM OOM OOM 0.8298 OOM 0.8722 OOM OOM OOM
COL Acc 0.7644 0.7328 0.7406 OOM 0.7556 0.8164 0.7696 OOM OOM OOM 0.7528 0.6988 0.8278 OOM 0.7688 OOM
F1 0.7454 0.7099 0.7169 OOM 0.7395 0.7973 0.7520 OOM OOM OOM 0.7114 0.6702 0.8046 OOM 0.7320 OOM
BIO PRO Acc 0.7286 0.7250 0.7367 0.7372 0.7277 0.7440 0.7286 0.7061 0.7475 0.7071 0.7322 0.7304 0.7232 0.7015 0.7367 0.7322
F1 0.6596 0.6432 0.6617 0.6645 0.6712 0.7058 0.6598 0.5838 0.6978 0.6026 0.6641 0.6653 0.6442 0.6386 0.6993 0.6647
$D$ Acc 0.7310 0.7233 0.7542 0.7488 0.7156 0.7802 0.7343 OOM 0.7776 0.7335 0.7632 0.7598 0.7632 OOM OOM OOM
F1 0.6676 0.6555 0.6707 0.7245 0.6384 0.7252 0.6687 OOM 0.7267 0.6402 0.6845 0.6841 0.6844 OOM OOM OOM
ENZ Acc 0.4650 0.4833 0.5183 0.5233 0.4800 0.4950 0.4783 0.4617 0.5267 0.4650 0.4417 0.4667 0.5250 0.4917 0.5317 0.4817
F1 0.4775 0.4781 0.5098 0.5204 0.4794 0.4237 0.4737 0.4592 0.5279 0.4628 0.4337 0.4695 0.5309 0.4902 0.5314 0.4772
HIV Acc 0.9696 0.9701 0.9695 0.9703 0.9687 0.9691 0.9696 0.9689 0.9728 0.9703 0.9686 0.9689 0.9701 0.9686 0.9705 0.9701
F1 0.3192 0.3466 0.2991 0.3622 0.2221 0.2861 0.3213 0.2537 0.3682 0.3413 0.2203 0.2518 0.2502 0.2225 0.3628 0.3448
TOX Acc 0.5548 0.5553 0.5529 0.5561 0.5523 0.5412 0.5514 0.5463 0.5561 0.5580 0.5283 0.5350 0.5384 0.5112 0.5565 0.5558
F1-i 0.9120 0.9114 0.9096 0.9135 0.9083 0.9109 0.9101 0.9121 0.9138 0.9137 0.8993 0.9004 0.9048 0.8986 0.9126 0.9123
F1-a 0.3628 0.3601 0.3418 0.3839 0.2240 0.3452 0.3765 0.3858 0.3850 0.3815 0.2020 0.2125 0.2180 0.1981 0.3847 0.3695
CHE MUT Acc 0.8041 0.8570 0.8199 0.8588 0.8094 0.8254 0.8091 0.8091 0.8409 0.8062 0.7643 0.7706 0.7790 0.7067 0.8607 0.8842
F1 0.8583 0.8907 0.8644 0.8909 0.8590 0.8642 0.8506 0.8558 0.8770 0.8512 0.8255 0.8296 0.8447 0.8012 0.8912 0.9070
NCI Acc 0.8158 0.8154 0.8146 0.8183 0.8169 0.7662 0.8160 0.7866 0.8187 0.7603 0.7816 0.7825 0.7852 0.6995 0.7560 0.8122
F1 0.8160 0.8153 0.8149 0.8184 0.8171 0.7708 0.8168 0.7914 0.8190 0.7673 0.7827 0.7831 0.7855 0.6997 0.7557 0.8119
BAC Acc 0.6711 0.6645 0.6477 0.7074 0.6756 0.6513 0.6749 0.7125 0.6803 0.6078 0.6019 0.6689 0.6294 0.6089 0.7099 0.6729
F1 0.7154 0.7351 0.7099 0.7342 0.7264 0.7366 0.7436 0.7400 0.7382 0.7055 0.7001 0.6545 0.7077 0.7047 0.7366 0.7186
PCB Acc 0.5456 0.5466 0.5451 0.5470 0.5461 0.5447 >3>3days OOM 0.5523 0.5414 OOM OOM OOM OOM OOM OOM
Mi-F1 0.9850 0.9853 0.9839 0.9853 0.9849 0.9836 >3>3days OOM 0.9860 0.9850 OOM OOM OOM OOM OOM OOM
Ma-F1 0.1200 0.1500 0.1317 0.1517 0.1333 0.0167 >3>3days OOM 0.2140 0.2064 OOM OOM OOM OOM OOM OOM
Table 5. Evaluation on graph regression. The best and second performance are highlighted in bold and underlined, respectively.
Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
MC 3-Cyc MAE 0.4396 0.3962 0.5119 0.4060 0.4290 0.4251 0.4238 0.0192 0.0022 0.0009 0.8776 0.4369 0.5410 0.4744 0.4231 0.3873
R2 0.6970 0.7546 0.8081 0.7493 0.7044 0.7107 0.5679 0.9996 0.9998 0.9999 0.1025 0.6850 0.5003 0.6062 0.7174 0.7595
4-Cyc MAE 0.2806 0.2536 0.5407 0.2509 0.2772 0.2720 0.2696 0.0148 0.0223 0.0062 0.6451 0.2745 0.5442 0.5398 0.2734 0.2204
R2 0.8229 0.8864 0.4011 0.8924 0.8331 0.8477 0.8406 0.9997 0.9990 0.9998 0.1424 0.8350 0.4032 0.4436 0.8606 0.8987
5-Cyc MAE 0.2782 0.1858 0.4608 0.2580 0.2399 0.2338 0.2659 0.0716 0.0340 0.0115 0.9021 0.2763 0.4532 0.4570 0.2180 0.1756
R2 0.8334 0.9364 0.4815 0.8866 0.8614 0.8936 0.8553 0.9862 0.9972 0.9996 0.0019 0.8190 0.6081 0.4939 0.9060 0.9408
6-Cyc MAE 0.3014 0.1895 0.4690 0.2841 0.2302 0.1788 0.2926 0.0862 0.0584 0.0368 0.8876 0.3035 0.4557 0.5027 0.1773 0.1784
R2 0.7926 0.9162 0.6189 0.8252 0.8803 0.9279 0.8073 0.9569 0.9956 0.9969 0.0021 0.7891 0.6219 0.6077 0.9301 0.9223
7-Cyc MAE 0.4011 0.2111 0.5898 0.2237 0.2852 0.1569 0.3942 0.1560 0.0563 0.0488 0.8273 0.4220 0.5713 0.5836 0.1500 0.2045
R2 0.5887 0.8635 0.4594 0.8469 0.7921 0.9354 0.6091 0.9460 0.9937 0.9946 0.0064 0.5487 0.4883 0.4692 0.9404 0.8770
8-Cyc MAE 0.4757 0.2632 0.5285 0.2841 0.2910 0.1377 0.4701 0.1153 0.0493 0.0399 0.7430 0.4790 0.5340 0.4808 0.1286 0.2592
R2 0.3570 0.7220 0.2212 0.7152 0.7553 0.9428 0.3840 0.9703 0.9949 0.9963 0.0330 0.3520 0.2051 0.3319 0.9466 0.7265
4-Path MAE 0.7154 0.4273 0.7340 0.2940 0.5273 0.1606 0.6361 0.0243 0.0150 0.0083 0.7784 0.5019 0.7316 0.7270 0.1512 0.4093
R2 0.1574 0.6349 0.0898 0.8619 0.5270 0.9459 0.3188 0.9921 0.9996 0.9998 0.0279 0.5390 0.0909 0.1310 0.9534 0.6579
5-Path MAE 0.6848 0.3949 0.6669 0.3628 0.5189 0.1559 0.6133 0.0134 0.0153 0.0092 0.7510 0.4846 0.6933 0.7053 0.1407 0.3805
R2 0.1050 0.6358 0.1675 0.6973 0.5015 0.9473 0.3254 0.9995 0.9995 0.9998 0.0283 0.5204 0.0823 0.0650 0.9562 0.6562
6-Path MAE 0.6164 0.3913 0.6594 0.3397 0.5629 0.1394 0.6930 0.0138 0.0133 0.0089 0.7571 0.4552 0.6627 0.6589 0.1299 0.3784
R2 0.2470 0.5958 0.1286 0.7429 0.4133 0.9539 0.1408 0.9995 0.9996 0.9998 0.0250 0.5459 0.1282 0.1280 0.9607 0.6076
4-Cliq MAE 0.3432 0.3449 0.3498 0.2504 0.3430 0.3427 0.3578 0.0090 0.0092 0.0005 0.3866 0.2355 0.3422 0.3424 0.1797 0.3401
R2 0.1612 0.1348 0.1086 0.8230 0.1213 0.1022 0.1057 0.9665 0.9964 0.9999 0.0890 0.8896 0.1604 0.1610 0.9002 0.1644
Tailed Tri MAE 0.3506 0.2886 0.3812 0.3470 0.3465 0.3665 0.3405 0.0189 0.0146 0.0034 0.8934 0.3791 0.3891 0.4095 0.3277 0.2747
R2 0.7884 0.8606 0.7303 0.7792 0.7813 0.7673 0.8053 0.9989 0.9995 0.9999 0.0176 0.7336 0.7008 0.6245 0.8177 0.8662
Chor. Cyc MAE 0.4377 0.3683 0.4390 0.3500 0.4216 0.3690 0.4246 0.0295 0.0336 0.0044 0.8720 0.4441 0.4413 0.4728 0.3671 0.3608
R2 0.6346 0.7109 0.6002 0.8121 0.6414 0.6848 0.6421 0.9970 0.9951 0.9999 0.0333 0.5746 0.5841 0.5684 0.7293 0.7149
Tri. Rec. MAE 0.4657 0.4034 0.4794 0.4410 0.4736 0.4068 0.4890 0.3619 0.3464 0.3271 0.8263 0.4286 0.4755 0.5409 0.4301 0.3943
R2 0.5211 0.6291 0.5335 0.5598 0.4921 0.6235 0.4702 0.6895 0.7063 0.7241 0.0338 0.5828 0.5037 0.3852 0.5691 0.6425
@@
Table 6. Efficiency evaluation. We report total training (Train (s)) and inference (Infer (s)) times with two decimal places, and GPU memory usage for training (T-Mem (MB)) and inference (I-Mem (MB)) in integers. “OOM” indicates out-of-memory.
Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
SN RED Train(s) 192.44 182.98 179.55 OOM 218.03 758.22 8365.60 OOM OOM OOM OOM OOM 371.86 OOM OOM OOM
Infer(s) 28.42 26.21 26.04 OOM 32.95 30.45 1953.91 OOM OOM OOM OOM OOM 79.35 OOM OOM OOM
T-Mem(MB) 287 288 310 OOM 287 2120 360 OOM OOM OOM OOM OOM 604 OOM OOM OOM
I-Mem(MB) 190 122 135 OOM 190 244 194 OOM OOM OOM OOM OOM 289 OOM OOM OOM
BIO ENZ Train(s) 26.05 26.34 25.23 72.33 31.95 38.85 375.25 312.46 350.25 470.05 120.48 67.50 75.29 495.23 57.53 125.45
Infer(s) 5.19 5.33 5.04 17.42 6.84 7.17 76.30 68.74 68.93 94.92 36.12 19.20 28.27 85.11 7.76 5.07
T-Mem(MB) 83 78 77 289 83 228 103 582 2369 1656 323 186 506 282 324 272
I-Mem(MB) 42 36 36 96 42 46 45 270 879 790 152 95 194 54 58 42
CHE MUT Train(s) 10.78 11.03 10.05 24.53 13.50 15.96 60.13 52.05 88.70 67.33 44.37 35.32 39.99 80.49 25.15 20.72
Infer(s) 3.86 3.89 3.51 7.94 5.27 4.01 17.74 18.91 21.18 25.10 15.80 11.63 15.18 29.78 7.64 3.94
T-Mem(MB) 42 44 44 152 42 82 46 422 775 1306 56 48 431 134 140 104
I-Mem(MB) 19 20 21 82 19 24 19 245 210 722 26 27 74 24 25 19
MC 4-Cyc Train(s) 166.50 160.95 140.88 458.70 180.57 1224.39 482.95 1180.51 645.97 1487.66 395.71 284.34 394.60 1770.94 300.21 304.95
Infer(s) 256.17 241.02 235.40 590.20 279.43 2058.62 1033.68 1776.91 920.13 2219.63 549.60 441.95 583.55 2675.22 461.87 262.40
T-Mem(MB) 88 85 85 228 88 380 104 1214 1734 3080 262 269 470 238 237 165
I-Mem(MB) 64 62 62 209 64 260 68 342 608 726 104 152 270 103 96 64